OpenStudy (anonymous):

Divide. Please show work. 28/-2

OpenStudy (anonymous):

\[\frac{28}{-2} = -\frac{28}{2} = -14\]

OpenStudy (anonymous):

thank you. Can you help me with -3/4*-18 Multiply. Show all work and be sure your answer is in simplest form.

OpenStudy (anonymous):

\[-\frac{3}{4} \times (-18) = -\frac{3}{4} \times (-\frac{72}{4})= \frac{216}{4} = 54\]